For Sunny Dale Bank the time required to serve customer at the drive by window is an important quality factor in competing with other banks in the city. (Show your computation)


a) Mean time to process a customer at the peak demand period is 5 minutes

b) Standard deviation of 1.5 minutes

c) Sample size of six customers

d) Design an X bar Chart that has type I error of 5% (z= 1.96)

e) After several weeks of sampling, two successive samples came in at 3.70 and 3.68 minutes respectively

f) Is the customer service process in statistical control?




1
Expert's answer
2021-10-07T15:10:03-0400

Mean time =5 min

Standard deviation = 1.5 minutes

n=6

type I error of 5 percent , z=1.96

UCL = 5.+1.96*1.5/(6)^0.5=6.2 min

LCL=5.-1.96*1.5/(6)^0.5=3.8 min

As the UCL and LCL contains the new successive samples of  3.70 and 3.68 minutes so the process is in control